Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him?

The sales price of a car is $12,590, which is 20% off the original price. What is the original price?
Solve the following equation for A : 2A/3 = 8 + 4A
hi im just studying for my S.A.T.'s and have difficulty with percentages.

Found 2 solutions by stanbon, keksjr:
Answer by stanbon(48545)

1).Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him?
Let original cost be "x"
Then x+0.35x=670
1.35x=670
x= $496.30
2).The sales price of a car is $12,590, which is 20% off the original price. What is the original price?
Let the original price be "x"
Then x-0.20x=12590
0.80x=12590
x=$15737.50

3). Solve the following equation for A : 2A/3 = 8 + 4A
Multiply thru by "3" to get:
2A = 24+12A
-10A=24
A=-12/5=-2.8
Cheers,
Stan H.

Answer by keksjr(7)

1).Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him?
Let original cost be "x"
Then x+0.35x=670
1.35x=670
x= $496.30